# Functional Determinants for Constrained Path Integrals in Minisuperspace Jackiw-Teitelboim Gravity

### v1

@@ 1 Introduction
The quantization of gravity remains one of the most fundamental open problems in theoretical physics. In contrast to the other fundamental interactions, general relativity resists straightforward quantization because of its diffeomorphism invariance, the absence of a fixed background structure, and its non-renormalizability when treated as an ordinary quantum field theory tHooft:1974toh. These difficulties make it challenging to construct a conventional quantum theory of the spacetime metric itself. Nevertheless, the path integral formulation offers a conceptually appealing framework for quantum gravity, in which transition amplitudes are expressed as functional integrals over equivalence classes of spacetime geometries Hawking:1978jz.
@@
where $S[g_{\mu\nu}]$ is the gravitational action and the integration is over all metric configurations interpolating between initial and final geometries ($g_{i}$ and $g_{f}$). The division by $\mathrm{Vol}(\mathrm{Diff})$ indicates the removal of diffeomorphism redundancy. For general relativity, $S[g_{\mu\nu}]$ is taken to be the Einstein-Hilbert action supplemented by the appropriate boundary term York:1972sj; Gibbons:1976ue, which, in units $8\pi G=1$, can be written as
@@
In the canonical formulation, the amplitude (1) may be viewed as formally generating solutions to the Wheeler-DeWitt equation DeWitt:1967ub, which plays the role of the fundamental equation for the wave functional of the universe. However, as it stands, the functional integral (1) over all possible Lorentzian geometries is ill-defined and extremely difficult to evaluate in a mathematically controlled way, especially in four spacetime dimensions.
These issues motivate the study of simplified models that more rigorously define the gravitational path integral while retaining essential physical properties. Lower-dimensional gravity theories have played an important role in this regard, clarifying conceptual and non-perturbative aspects of quantum gravity. In particular, two-dimensional dilaton gravity captures key features of gravitational dynamics such as black hole thermodynamics, holographic AdS/CFT dualities, and non-perturbative effects while often remaining exactly solvable or useful to the approximation. Among these theories, Jackiw-Teitelboim gravity Jackiw:1984je; Teitelboim:1983ux has attracted considerable attention. Most work has focused on Jackiw-Teitelboim gravity in anti-de Sitter spacetime Almheiri:2014cka; Jensen:2016pah; Maldacena:2016upp; Engelsoy:2016xyb; Stanford:2017thb; Saad:2019lba, but there has also been significant progress in developing its de Sitter counterpart Maldacena:2019cbz; Stanford:2019vob; Cotler:2019nbi. The de Sitter version is particularly relevant for quantum cosmology, given the observed positive cosmological constant, and thus provides a natural testing ground for proposals about cosmological initial conditions and gravitational path integrals.
@@
where $\Lambda$ is the cosmological constant, $\gamma$ is the induced metric on the boundary, and $\mathcal{K}$ is the corresponding extrinsic curvature. The boundary term ensures a well-posed variational principle for fixed boundary conditions. The dilaton field $\phi$ appears linearly and therefore acts as a Lagrange multiplier enforcing the constant-curvature constraint $R=2\Lambda$. This constraint structure is precisely what makes Jackiw-Teitelboim gravity an ideal model for studying the structure of the gravitational path integral. Under asymptotically AdS2 boundary conditions, it is well established that the bulk path integral reduces to the boundary Schwarzian theory Maldacena:2016upp; Engelsoy:2016xyb; Stanford:2017thb; Saad:2019lba. Moreover, because the path integral over the dilaton $\phi$ can be carried out exactly, the remaining metric path integral is localized onto constant-curvature geometries, yielding a purely geometric representation of the Jackiw-Teitelboim path integral. In particular, integrating out $\phi$ and then evaluating the resulting constrained path integral, we obtain the following formula,
@@
where $\delta[\cdots]$ denotes the functional delta, and $\phi_{\rm bdy}$ collectively denotes the dilaton values on $\partial\mathcal{M}$ (in particular the initial and final values $\phi_{i}$ and $\phi_{f}$ in the present amplitude). In the last step, we have used the functional generalization of Eq. (14) to evaluate the $\delta$-functional by expanding around the classical metric $g_{\text{cl}}$ satisfying $R(g_{\rm cl})=2\Lambda$ with the prescribed boundary conditions. The delta functional in Eq. (4) must be carefully calculated and requires computing the determinant of the linearised constraint operator. Once the determinant is correctly computed, Eq. (4) provides the gravitational transition amplitude for any Lorentzian geometry in Jackiw-Teitelboim gravity. In the geometric formulation, the boundary Schwarzian mode is encoded in this determinant. Although evaluating the determinant for fluctuations of the full metric $g_{\mu\nu}$ (with an appropriate gauge fixing) is technically complicated, it is essential for verifying consistency with the Schwarzian description. We will provide a detailed calculation and analysis of this determinant in a separate publication.
@@
Instead, in this paper, we evaluate the constrained path integral in the minisuperspace approximation, which provides a practical approach for evaluating the path integral of quantum gravity. This minisuperspace approximation truncates the infinite degrees of freedom of the metric to a finite variable by assuming a high degree of symmetry (e.g., homogeneity and isotropy). In our previous collaborative work Honda:2024hdr, we evaluated the minisuperspace path integral for Jackiw-Teitelboim gravity by performing the path integrals over the dilaton and the scale factor separately. This allowed us to evaluate the Jackiw-Teitelboim path integral without resorting to the constrained path integral of Eq. (4). In this paper, we first derive and evaluate the constrained path integral in minisuperspace and demonstrate that it yields the same results as those reported in Honda:2024hdr. Furthermore, to determine the prefactor in the path integral, we compute the functional determinant using the Gelfand-Yaglom theorem Gelfand:1959nq. This theorem provides a powerful method for computing determinants of one-dimensional differential operators by relating them directly to solutions of the corresponding classical equations of motion Dunne:2007rt; Coleman:1985rnk. Thus, this work provides a full evaluation of the Lorentzian path integral in Jackiw-Teitelboim gravity. Then, we generalize our analysis to Jackiw-Teitelboim gravity with dilaton potential $U(\phi)=m^{2}\phi^{2}$, which includes the pure Jackiw-Teitelboim case as $m^{2}=0$. As a further application of our methods, we discuss the calculation for Bianchi IX quantum cosmology Hawking:1984wn; Wright:1984wm; Amsterdamski:1985qu; Jensen:1990xc; DiazDorronsoro:2018wro; Feldbrugge:2018gin; Janssen:2019sex; Lehners:2024kus, which involves several minisuperspace variables and exhibits a more complex constraint structure. This demonstrates the broader applicability of functional determinant techniques in minisuperspace quantum cosmology and quantum gravity.
@@
The paper is organized as follows. In Section 2, we review the minisuperspace reduction of Jackiw-Teitelboim gravity and rewrite the minisuperspace path integral as a constrained path integral by integrating out the dilaton. In Section 3, we present the general treatment of functional delta constraints, and compute the relevant functional determinant using the Gelfand-Yaglom theorem. In Section 4, we perform the lapse integral over $N>0$ to obtain the Lorentzian transition amplitude of Jackiw-Teitelboim gravity, and review the results of Honda:2024hdr relevant for our discussion. In Section 5, we extend the analysis to general Jackiw-Teitelboim gravity with a dilaton potential $U(\phi)=m^{2}\phi^{2}$. In Section 6, we apply our analysis to Bianchi IX quantum cosmology and compute the functional determinant. Section 7 concludes with a summary and future directions.
@@ 2 Jackiw-Teitelboim Gravity
The Jackiw-Teitelboim gravity provides the simplest example of two-dimensional dilaton gravity. We consider the minisuperspace reduction to one spatial dimension with metric $\mathrm{d}s^{2}=-N(t)^{2}\mathrm{d}t^{2}+a(t)^{2}\mathrm{d}x^{2}$, where $x$ has period $2\pi\ell$ with $\ell$ the spatial coordinate length and $N(t)$ is the lapse function. The minisuperspace Jackiw-Teitelboim action takes the form Honda:2024hdr
@@
and we transformed the lapse function $N\to N/a\phi$ (see Ref. Honda:2024hdr for the details). We impose Dirichlet boundary conditions on both fields,
@@
These specify the initial and final geometric configurations, appropriate for computing transition amplitudes in quantum cosmology. Once the boundary conditions are imposed, we can evaluate the gravitational path integral. Through the Batalin-Fradkin-Vilkovisky (BFV) formalism Fradkin:1975cq; Batalin:1977pb, the gravitational path integral (1) of Jackiw-Teitelboim gravity reduces to a quantum mechanical-like path integral Halliwell:1988wc; Halliwell:1988ik
@@
which is the integral over the proper time $N\left(t_{1}-t_{0}\right)$ between the initial and final configurations. Much of the recent work in Lorentzian quantum cosmology Feldbrugge:2017kzv; Lehners:2023yrj has focused on giving a rigorous definition of the lapse integral $\int\mathrm{d}N$ using Picard-Lefschetz theory Witten:2010cx. This is crucial for resolving ambiguities between different cosmological proposals, such as the Hartle-Hawking no-boundary proposal Hartle:1983ai and Vilenkin’s tunneling proposal Vilenkin:1984wp, as well as related early proposals Linde:1983cm; Linde:1984ir; Rubakov:1984bh; Zeldovich:1984vk. However, to address this problem completely, we must combine resurgence theory with a systematic analysis of Lefschetz thimbles Honda:2024aro. Within this framework, the minisuperspace path integral can be evaluated consistently, and the ambiguity in the wave function of the universe can be resolved.
By performing the path integrals over the dilaton and the scale factor separately, we obtain the following propagator Honda:2024hdr,
@@
When we have integrated the propagator for $N$ over all positive values of the lapse function, we can get the Lorentzian transition amplitude of Jackiw-Teitelboim gravity. The detailed discussion is given by Ref. Fanaras:2021awm; Honda:2024hdr, and we provide a brief review in Section 4.
@@
A central technical difficulty in evaluating such constrained path integrals is the proper treatment of the functional delta and its associated Jacobian. Performing a functional delta function gives rise to functional determinants of differential operators, which must be carefully calculated. The Gelfand-Yaglom theorem Gelfand:1959nq provides an elegant way to compute these determinants: for a differential operator defined on a finite interval with specified boundary conditions, the theorem expresses its functional determinant in terms of the boundary value of the solution to an associated initial-value problem. This method has been widely used in quantum field theory: detailed reviews can be found in Refs. Coleman:1985rnk; Dunne:2007rt. In the following, we use the Gelfand-Yaglom theorem to analyze the constrained path integral.
@@ 3.2 Gelfand-Yaglom theorem
The Gelfand-Yaglom theorem Gelfand:1959nq states that for a one-dimensional second-order differential operator of the form
@@
which coincides with the result (9) obtained in Ref. Honda:2024hdr. Thus, we have confirmed that our evaluation of the constrained path integral using the Gelfand-Yaglom theorem to compute the functional determinant is justified.
@@ 4 Lorentzian Transition Amplitude and Saddle-Point Structure
Having obtained the propagator at fixed lapse $N$, we construct the full Lorentzian transition amplitude by integrating over the positive lapse function, $N>0$. This lapse integral exhibits a nontrivial saddle-point structure, whose contributing saddles and associated steepest-descent contours encode characteristic features of Lorentzian quantum cosmology Feldbrugge:2017kzv; Honda:2024aro. A detailed analysis in the context of Jackiw-Teitelboim gravity is presented in Ref. Fanaras:2021awm; Honda:2024hdr; here we summarize the results relevant for our discussion.
@@
To understand the saddle-point structure more clearly, we use Lefschetz thimble techniques Witten:2010zr; Cristoforetti:2012su. Changing variables to $N=e^{x}$ with $x\in\mathbb{C}$, the exponent becomes
@@
which lies along the imaginary axis (${\rm Im}N<0$). This represents a Euclidean saddle point in the complex lapse plane so that the path integral localizes onto a smooth complex geometry realizing the Lorentzian continuation of Euclidean evolution. This precisely captures the picture of Hartle-Hawking no-boundary proposal Hartle:1983ai.
@@ 6 Bianchi IX Quantum Cosmology
To demonstrate the broader applicability of our methods, we consider the application to Bianchi IX quantum cosmology. This model contains several independent degrees of freedom and provides a standard approach for studying anisotropic dynamics in quantum cosmology Hawking:1984wn; Wright:1984wm; Amsterdamski:1985qu; Jensen:1990xc; DiazDorronsoro:2018wro; Feldbrugge:2018gin; Janssen:2019sex; Lehners:2024kus. The Bianchi IX model describes a homogeneous but anisotropic spacetime with spatial topology $S^{3}$. In the biaxial set-up, the metric can be parametrized by two time-dependent scale factors, which we denote by $p(t)$ and $q(t)$: roughly speaking, $p$ controls the overall volume while $q/p$ measures the anisotropy. The corresponding minisuperspace action reads DiazDorronsoro:2018wro

### v2

@@ 1 Introduction
The quantization of gravity remains one of the most fundamental open problems in theoretical physics. In contrast to the other fundamental interactions, general relativity resists straightforward quantization because of its diffeomorphism invariance, the absence of a fixed background structure, and its non-renormalizability when treated as an ordinary quantum field theory ’t Hooft and Veltman (1974). These difficulties make it challenging to construct a conventional quantum theory of the spacetime metric itself. Nevertheless, the path integral formulation offers a conceptually appealing framework for quantum gravity, in which transition amplitudes are expressed as functional integrals over equivalence classes of spacetime geometries Hawking (1978).
@@
where $S[g_{\mu\nu}]$ is the gravitational action and the integration is over all metric configurations interpolating between initial and final geometries ($g_{i}$ and $g_{f}$). The division by $\mathrm{Vol}(\mathrm{Diff})$ indicates the removal of diffeomorphism redundancy. For general relativity, $S[g_{\mu\nu}]$ is taken to be the Einstein-Hilbert action supplemented by the appropriate boundary term York (1972); Gibbons and Hawking (1977), which, in units $8\pi G=1$, can be written as
@@
In the canonical formulation, the amplitude (1) may be viewed as formally generating solutions to the Wheeler-DeWitt equation DeWitt (1967), which plays the role of the fundamental equation for the wave functional of the universe. However, as it stands, the functional integral (1) over all possible Lorentzian geometries is ill-defined and extremely difficult to evaluate in a mathematically controlled way, especially in four spacetime dimensions.
These issues motivate the study of simplified models that more rigorously define the gravitational path integral while retaining essential physical properties. Lower-dimensional gravity theories have played an important role in this regard, clarifying conceptual and non-perturbative aspects of quantum gravity. In particular, two-dimensional dilaton gravity captures key features of gravitational dynamics such as black hole thermodynamics, holographic AdS/CFT dualities, and non-perturbative effects while often remaining exactly solvable or useful to the approximation. Among these theories, Jackiw-Teitelboim gravity Jackiw (1985); Teitelboim (1983) has attracted considerable attention. Most work has focused on Jackiw-Teitelboim gravity in anti-de Sitter spacetime Almheiri and Polchinski (2015); Jensen (2016); Maldacena et al. (2016); Engelsöy et al. (2016); Stanford and Witten (2017); Saad et al. (2019), but there has also been significant progress in developing its de Sitter counterpart Maldacena et al. (2021); Stanford and Witten (2020); Cotler et al. (2020). The de Sitter version is particularly relevant for quantum cosmology, given the observed positive cosmological constant, and thus provides a natural testing ground for proposals about cosmological initial conditions and gravitational path integrals.
@@
where $\Lambda$ is the cosmological constant, $\gamma$ is the induced metric on the boundary, and $\mathcal{K}$ is the corresponding extrinsic curvature. The boundary term ensures a well-posed variational principle for fixed boundary conditions. The dilaton field $\phi$ appears linearly and therefore acts as a Lagrange multiplier enforcing the constant-curvature constraint $R=2\Lambda$. This constraint structure is precisely what makes Jackiw-Teitelboim gravity an ideal model for studying the structure of the gravitational path integral. Under asymptotically AdS2 boundary conditions, it is well established that the path integral reduces to the boundary Schwarzian theory Maldacena et al. (2016); Engelsöy et al. (2016); Stanford and Witten (2017); Saad et al. (2019). Moreover, because the path integral over the dilaton $\phi$ can be carried out exactly, the remaining metric path integral is localized onto constant-curvature geometries, yielding a purely geometric representation of the Jackiw-Teitelboim path integral. In particular, integrating out $\phi$ and then evaluating the resulting constrained path integral, we obtain the following formula,
@@
where $\delta[\cdots]$ denotes the functional delta, and $\phi_{\rm bdy}$ collectively denotes the dilaton values on $\partial\mathcal{M}$ (in particular the initial and final values $\phi_{i}$ and $\phi_{f}$ in the present amplitude). In the last step, we have used the functional generalization of Eq. (14) to evaluate the $\delta$-functional by expanding around the classical metric $g_{\text{cl}}$ satisfying $R(g_{\rm cl})=2\Lambda$ with the boundary conditions. The delta functional in Eq. (4) must be carefully calculated and requires computing the determinant of the linearised constraint operator. Once the determinant has been evaluated, Eq. (4) yields the gravitational transition amplitude for an arbitrary Lorentzian geometry in Jackiw-Teitelboim gravity. In practice, however, computing this determinant for fluctuations of the full metric $g_{\mu\nu}$ with an appropriate gauge fixing is technically challenging. A complete derivation and detailed analysis will be presented in a separate publication.
@@
Instead, in this paper, we evaluate the constrained path integral in the minisuperspace approximation, which provides a practical approach for evaluating the path integral of quantum gravity. This minisuperspace approximation truncates the infinite degrees of freedom of the metric to a finite variable by assuming a high degree of symmetry (e.g., homogeneity and isotropy). In our previous collaborative work Honda et al. (2025), we evaluated the minisuperspace path integral for Jackiw-Teitelboim gravity by performing the path integrals over the dilaton and the scale factor separately. This allowed us to evaluate the Jackiw-Teitelboim path integral without resorting to the constrained path integral of Eq. (4). In this paper, we first derive and evaluate the constrained path integral in minisuperspace and demonstrate that it yields the same results as those reported in Honda et al. (2025). Furthermore, to determine the prefactor in the path integral, we compute the functional determinant using the Gelfand-Yaglom theorem Gelfand and Yaglom (1960). This theorem provides a powerful method for computing determinants of one-dimensional differential operators by relating them directly to solutions of the corresponding classical equations of motion Coleman (1985); Kirsten and McKane (2004); Dunne (2008). Thus, this work provides a full evaluation of the Lorentzian path integral in Jackiw-Teitelboim gravity. Then, we generalize our analysis to Jackiw-Teitelboim gravity with dilaton potential $U(\phi)=m^{2}\phi^{2}$, which includes the pure Jackiw-Teitelboim case as $m^{2}=0$. As a further application of our methods, we discuss the calculation for Bianchi IX quantum cosmology Hawking and Luttrell (1984); Wright and Moss (1985); Amsterdamski (1985); Jensen et al. (1991); Diaz Dorronsoro et al. (2018); Feldbrugge et al. (2018); Janssen et al. (2019); Lehners (2025); Lo Franco and Montani (2026), which involves several minisuperspace variables and exhibits a more complex constraint structure. This demonstrates the broader applicability of functional determinant techniques in minisuperspace quantum cosmology and quantum gravity.
@@
The paper is organized as follows. In Section 2, we review the minisuperspace reduction of Jackiw-Teitelboim gravity and rewrite the minisuperspace path integral as a constrained path integral by integrating out the dilaton. In Section 3, we present the general treatment of functional delta constraints, and compute the relevant functional determinant using the Gelfand-Yaglom theorem. In Section 4, we perform the lapse integral over $N>0$ to obtain the Lorentzian transition amplitude of Jackiw-Teitelboim gravity, and review the results of Honda et al. (2025) relevant for our discussion. In Section 5, we extend the analysis to general Jackiw-Teitelboim gravity with a dilaton potential $U(\phi)=m^{2}\phi^{2}$. In Section 6, we apply our analysis to Bianchi IX quantum cosmology and compute the functional determinant. Section 7 concludes with a summary and future directions.
@@ 2 Jackiw-Teitelboim Gravity
The Jackiw-Teitelboim gravity provides the simplest example of two-dimensional dilaton gravity. We consider the minisuperspace reduction to one spatial dimension with metric $\mathrm{d}s^{2}=-N(t)^{2}\mathrm{d}t^{2}+a(t)^{2}\mathrm{d}x^{2}$, where $x$ has period $2\pi\ell$ with $\ell$ the spatial coordinate length and $N(t)$ is the lapse function. The minisuperspace Jackiw-Teitelboim action takes the form Honda et al. (2025)
@@
and we transformed the lapse function $N\to N/a\phi$ (see Ref. Honda et al. (2025) for the details). We impose Dirichlet boundary conditions on both fields,
@@
These specify the initial and final geometric configurations, appropriate for computing transition amplitudes in quantum cosmology. Once the boundary conditions are imposed, we can evaluate the gravitational path integral. Through the Batalin-Fradkin-Vilkovisky (BFV) formalism Fradkin and Vilkovisky (1975); Batalin and Vilkovisky (1977), the gravitational path integral (1) of Jackiw-Teitelboim gravity reduces to a quantum mechanical-like path integral Halliwell (1988); Halliwell and Louko (1989)
@@
which is the integral over the proper time $N\left(t_{1}-t_{0}\right)$ between the initial and final configurations. Much of the recent work in Lorentzian quantum cosmology Feldbrugge et al. (2017); Lehners (2023) has focused on giving a rigorous definition of the lapse integral $\int\mathrm{d}N$ using Picard-Lefschetz theory Witten (2011). This is crucial for resolving ambiguities between different cosmological proposals, such as the Hartle-Hawking no-boundary proposal Hartle and Hawking (1983) and Vilenkin’s tunneling proposal Vilenkin (1984), as well as related early proposals Linde (1984a, b); Rubakov (1984); Zeldovich and Starobinsky (1984). However, to address this problem completely, we must combine resurgence theory with a systematic analysis of Lefschetz thimbles Honda et al. (2024). Within this framework, the minisuperspace path integral can be evaluated consistently, and the ambiguity in the wave function of the universe can be resolved.
By performing the path integrals over the dilaton and the scale factor separately, we obtain the following propagator Honda et al. (2025),
@@
When we have integrated the propagator for $N$ over all positive values of the lapse function, we can get the Lorentzian transition amplitude of Jackiw-Teitelboim gravity. The detailed discussion is given by Ref. Fanaras and Vilenkin (2022); Honda et al. (2025), and we provide a brief review in Section 4.
@@
A central technical difficulty in evaluating such constrained path integrals is the proper treatment of the functional delta and its associated Jacobian. Performing a functional delta function gives rise to functional determinants of differential operators, which must be carefully calculated. The Gelfand-Yaglom theorem Gelfand and Yaglom (1960) provides an elegant way to compute these determinants: for a differential operator defined on a finite interval with specified boundary conditions, the theorem expresses its functional determinant in terms of the boundary value of the solution to an associated initial-value problem. This method has been widely used in quantum field theory: detailed reviews can be found in Refs. Coleman (1985); Dunne (2008). In the following, we use the Gelfand-Yaglom theorem to analyze the constrained path integral.  [^2] Another study Ailiga et al. (2025) employs the Gelfand-Yaglom method to compute the functional determinants in Lorentzian quantum cosmology.
@@ 3.2 Gelfand-Yaglom theorem
The Gelfand-Yaglom theorem Gelfand and Yaglom (1960) states that for a one-dimensional second-order differential operator of the form
@@
which coincides with the result (9) obtained in Ref. Honda et al. (2025). Thus, we have confirmed that our evaluation of the constrained path integral using the Gelfand-Yaglom theorem to compute the functional determinant is justified.
@@ 4 Lorentzian Transition Amplitude and Saddle-Point Structure
Having obtained the propagator at fixed lapse $N$, we construct the full Lorentzian transition amplitude by integrating over the positive lapse function, $N>0$. This lapse integral exhibits a nontrivial saddle-point structure, whose contributing saddles and associated steepest-descent contours encode characteristic features of Lorentzian quantum cosmology Feldbrugge et al. (2017); Honda et al. (2024). A detailed analysis in the context of Jackiw-Teitelboim gravity is presented in Ref. Fanaras and Vilenkin (2022); Honda et al. (2025); here we summarize the results relevant for our discussion.
@@
To understand the saddle-point structure more clearly, we use Lefschetz thimble techniques Witten (2010); Cristoforetti et al. (2012). Changing variables to $N=e^{x}$ with $x\in\mathbb{C}$, the exponent becomes
@@
which lies along the imaginary axis (${\rm Im}N<0$). This represents a Euclidean saddle point in the complex lapse plane so that the path integral localizes onto a smooth complex geometry realizing the Lorentzian continuation of Euclidean evolution. This precisely captures the picture of Hartle-Hawking no-boundary proposal Hartle and Hawking (1983).
@@ 6 Bianchi IX Quantum Cosmology
To demonstrate the broader applicability of our methods, we consider the application to Bianchi IX quantum cosmology. This model contains several independent degrees of freedom and provides a standard approach for studying anisotropic dynamics in quantum cosmology Hawking and Luttrell (1984); Wright and Moss (1985); Amsterdamski (1985); Jensen et al. (1991); Diaz Dorronsoro et al. (2018); Feldbrugge et al. (2018); Janssen et al. (2019); Lehners (2025); Lo Franco and Montani (2026). The Bianchi IX model describes a homogeneous but anisotropic spacetime with spatial topology $S^{3}$. In the biaxial set-up, the metric can be parametrized by two time-dependent scale factors, which we denote by $p(t)$ and $q(t)$: roughly speaking, $p$ controls the overall volume while $q/p$ measures the anisotropy. The corresponding minisuperspace action reads Diaz Dorronsoro et al. (2018)